OpenStudy (anonymous):

Assume that x and y are both differentiable functions of t and find the required values of dy/dt and dx/dt. x^2 + y^2 = 100 (a) Find dy/dt, when x = 6, y = 8 given that dx/dt = 4. (b) Find dx/dt, when x = 8, y = 6, given that dy/dt = −2.

OpenStudy (anonymous):

What do I do?

OpenStudy (anonymous):

differentiate with respect to t

OpenStudy (anonymous):

\[\frac{ d }{ dt }\left( x^2 \right)=2 x \frac{ dx }{ dt }\]

OpenStudy (anonymous):

So I just have to plug in the 2(6)+8^2?

OpenStudy (anonymous):

\[2 x \frac{ dx }{ dt }+2 y \frac{ dy }{ dt }=0,x \frac{ dx }{ dt }+y \frac{ dy }{ dt }=0\]

OpenStudy (anonymous):

but they are giving me these values of x=6 and y=8... and then values of x=8 and y=6..

OpenStudy (anonymous):

SOMEBODY please help!!???

OpenStudy (phi):

surji wrote out most of the solution for you. you start with x^2 + y^2 = 100 you differentiate both sides (and all terms) with respect to t as posted above you get 2 x dx/dt + 2y dy/dt =0 or (dividing by 2) x dx/dt + y dy/dt =0 now replace the variables with the given values and evaluate.

OpenStudy (anonymous):

So... is it 2(6)+2(8)=0??? and 2(8)+2(6)??

OpenStudy (anonymous):

I'm taking this course online and I have no idea what I'm doing

OpenStudy (phi):

(a) Find dy/dt, when x = 6, y = 8 given that dx/dt = 4. they give you 3 out of the 4 unknowns in: x dx/dt + y dy/dt =0 dy/dt is the 4th unknown that you have to solve for.

OpenStudy (anonymous):

huh? sooooooo..... what does it mean when x=6 and y=8 and dx/dt=4? and what do I do here... I am lost to this stuff since I have no person to ask and walk me through the steps

OpenStudy (anonymous):

I get part of it =0 but I don't understand where to plug in what and when..

OpenStudy (phi):

It means replace x with 6, replace y with 8 and dx/dt with 4 in the equation x dx/dt + y dy/dt =0 what do you get ?

OpenStudy (anonymous):

6^2+8^2=100 it says it =4 on the next part, how do they get 4?

OpenStudy (anonymous):

2(6)=12 not 4

OpenStudy (anonymous):

if the function changes to 2x

OpenStudy (phi):

Let's start at the beginning. they give you an equation x^2 + y^2 = 100 which is the equation of a circle with center at (0,0) and radius = 10 they tell you both x and y are "functions of t" in other words, where you are on the circle is determined by time. Think of a satellite in orbit, or a spot on a wheel that is turning.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i picture an x y axis here right?

OpenStudy (phi):

as you move along the circle, the velocity in the x direction (denoted as dx/dt ) changes. for example, at the top of the circle you are moving mostly left/right at the side of the circle, you are moving mostly up/down. the idea is the velocity in the x direction (dx/dt) and the velocity in the y direction (dy/dt) change depending on where you are

OpenStudy (phi):



OpenStudy (phi):

using implicit differentiation we can find an equation that tells us some info about the velocity dx/dt and dy/dt Hopefully you know how to find \[ x \frac{dx}{dt} + y \frac{dy}{dt} = 0 \] if we are at the point (6,8), and are moving in the x direction at speed 4, how fast are we moving in the y direction?

OpenStudy (phi):